3. Find the equation of the line parallel to 5x + 6y – 10 = 0 at a distance 61 from (-3,7).

4. Find the equation of the line parallel to 3x + 4y = 20 and at a distance of 5 units from this

line.


1
Expert's answer
2021-10-18T15:34:40-0400

3. The equation of the line parallel to "5x + 6y \u2013 10 = 0" is


"5x + 6y +b = 0"

Distance from point "(-3,7)" to the line is


"d=\dfrac{|5(-3)+6(7)+b|}{\sqrt{5^2+6^2}}=61"

"|b-27|=61\sqrt{61}"

"b=27+61\sqrt{61}"

The equation of the line is


"5x + 6y +27+61\sqrt{61} = 0"



"b=27-61\sqrt{61}"

The equation of the line is


"5x + 6y +27-61\sqrt{61} = 0"




4. The equation of the line parallel to "3x + 4y = 20" is


"3x + 4y = b"


Take the point "(4, 2)"


"3(4) + 4(2) = 20, True"

Distance from point "(4, 2)" to the line "3x + 4y = b" is


"d=\dfrac{|3(4)+4(2)-b|}{\sqrt{3^2+4^2}}=5"

"|20-b|=25"

"b=-5"

The equation of the line parallel to "3x + 4y = 20" is


"3x + 4y = -5"




"b=45"

The equation of the line parallel to "3x + 4y = 20" is


"3x + 4y = 45"